44The following random sample of 28 female basketball player heights, in inches, is:

63 71 69 65 73 84 70 69 67 74 75 68 65 63 67 69 68 72 73 75 72 75 73 68 69 74 65 65
(Σx = 1961       Σx2  = 137,911)

The shape of the box plot representing this distribution of female basketball player heights is:
Answer #1

Answer:-  

Arrange the data into ascending order first:-

63,63,65,65,65,65,67,67,68,68,68,69,69,69,69,70,71,72,72,73,73,73,74,74,75,75,75,84

n=28 Ź Median nus que therefore median (h+)/2 = (28+1)/2 = 12:5 is not ari integer value, the median is average of 14th and 1

los pos 103) = (hti) 25 (28+) 7 - 21.75 Since 21.75 is not integer number Osus computed by interpolating blw the value pocateप 80 8 न hany Median 69 Max:04 mins63 ७. -67 2013 TOR-6 Outlier 84 नप ती तेण 67 66 64 G